A manufacturer claims that only 4% of his products supplied by him are defective. A random sample of 600 products contains 36 defectives. Test the claim of the manufacturer.


1
Expert's answer
2022-05-15T09:31:33-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\le 0.04"

"H_a:p>0.04"

This corresponds to a right-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a right-tailed test is "z_c = 1.6449."

The rejection region for this right-tailed test is "R = \{z: z > 1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{N}}}=\dfrac{\dfrac{36}{600}-0.04}{\sqrt{\dfrac{0.04(1-0.04)}{600}}}=2.5"

Since it is observed that "z = 2.5 > 1.6449=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p=P(Z>2.5)=0.00621," and since "p=0.00621<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population proportion "p" is greater than 0.04, at the "\alpha = 0.05" significance level.